Discuss with colleagues and IITians

A CAR HAVING A MASS OF 1000Kg DEFLECTS ITS SPRING 4 cm UNDER ITS LOAD.DETERMINE THE NATURAL FREQUENCY OF THE CAR IN VERTICAL DIRECTION.

TEJASVI BHASKAR
14 Years agoGrade

1 Answer

Chetan Mandayam Nayakar
ApprovedApproved Tutor Answer14 Years ago

F=kx,k=F/x=9800/0.04

m=1000kg

ω=√(k/m)